if the length of the rectangular label is 12 centimeters, what is the radius of the base of the can

Answer by stanbon(75887)

if the length of the rectangular label is 12 centimeters, what is the radius of the base of the can
----
circumference = 2*pi*r
----
Equation:
2(pi)r = 12
r = 12/(2*pi) = 6/pi cm
============================
Cheers,
Stan H.